# Wave-packet dynamics in pseudo-Hermitian lattices: Coexistence of Hermitian and non-Hermitian wavefronts

### v1

@@ III.3 Effects of weak disorder
In this section, we investigate the effect of weak disorder, which does not localize the wave-packet on the scale of the system size. While such disorder barely alters the dynamics in Hermitian systems, we will show that in non-Hermitian systems, it can lead to the emergence of a wave-packet, similar to the effect seen earlier (caused by the reflection of the Hermitian wave-packet from the wall). However, in this case, the cause is the disorder. As we will discuss in detail below, the main reason for this behavior is that in non-Hermitian systems, the tail of the wave-packet in position space (typically insignificant in Hermitian systems) dominates the long-time dynamics due to the transformation in Eq. (6). From another perspective, weak disorder causes weak scattering to all $k$-states, but even small occupations of certain $k$-states will grow exponentially in non-Hermitian systems. This phenomenon occurs only when the initial width of the wave-packet exceeds a certain threshold, denoted as $\sigma_{c,\mathrm{dis}}$, which depends on the disorder strength and differs from the previously discussed critical value $\sigma_{c,\mathrm{ref}}$.
@@
Examining the wave-packet in $k$-space, the disorder enables transitions from the center of the wave-packet (initially positioned at $k_{0}$) to all other $k$ values. In the case of a Hermitian Hamiltonian, this scattering process continues until it reaches a saturation value determined by the disorder strength $w$ and the system size. However, since the Hamiltonian is non-Hermitian, each $k$-value also experiences exponential growth or decay. In particular, the most significant $k$-value which is $k=-\pi/2$ (since it exhibits the fastest growth, as was discussed earlier), will eventually surpass the initial Gaussian wave-packet in magnitude. This explains the transition presented in the figure. To estimate the transition time, we can develop approximations for the amplitudes of both the initial Gaussian wave-packet and the $k=-\pi/2$ packet. As we saw before, up to intermediate times the Gaussian can be well described by the continuum limit. We thus find that
@@
where $k_{m}$ is the maximum of the wave-packet, and $b=\ln\left(\frac{t_{r}}{t_{l}}\right)\sqrt{t_{l}t_{r}}$. On the other hand, using perturbation theory and averaging over the disorder distribution, we find that the intermediate saturation value and the corresponding time are
$V_{s}\approx\frac{8}{3}\frac{w^{2}}{N}\frac{1}{(t_{r}+t_{l})^{2}},\,t_{s}\approx\frac{\pi\sqrt{2}}{(t_{r}+t_{l})^{2}}.$ (27)
Adding the exponential growth of $k=-\frac{\pi}{2}$, we get that
$|\psi(k=-\frac{\pi}{2},t)|^{2}\approx V_{s}e^{2(t_{l}-t_{r})(t-t_{s})}.$ (28)
We have verified that these approximations are in a good agreement with the numerical results (see Appendix E. for more details). By comparing Eq. (26) and (28), we obtain a quadratic equation, and its solution readily gives the transition time, which is marked in green on the figure. Finally, it is worth noting that even if no disorder is considered ($w=0$), numerical artifacts may induce such transition. This is because the numerical accuracy of the computer may not be sufficient to capture the evolution of the tail of the wave-packet, effectively introducing a disorder term with amplitude related to the computer precision threshold. To overcome this (for example, to plot the disorder-less scenario that was presented in Fig. 4) one may need to increase the precision by using an arbitrary precision library (such as Advanpix [68] in Matlab).
@@ Appendix D Calculation of the transition time of the non-Hermitian reflection
Keeping $a_{0}$ to be some reference lattice constant (where $a\ll a_{0}$), we have $v_{h}=2a\sqrt{t_{r}t_{l}}$ where $\sqrt{t_{l}t_{r}}=\frac{a_{0}^{2}}{a^{2}}\sqrt{t_{l,0}t_{r,0}},$ and $r=\left(\frac{t_{r,0}}{t_{l,0}}\right)^{a/2a_{0}}$ (see Appendix C). Plugging this into Eq. (80,81), we finally obtain

### v2

@@ III.3 Effects of weak disorder
In this section, we investigate the effect of weak disorder, which does not localize the wave-packet on the scale of the system size. Related work on “non-Hermitian jumps” considers strong disorder which leads to Anderson localization, and thus addresses a different parameter regime [71, 72, 73, 74, 75, 76, 77, 78]. While weak disorder barely alters the dynamics in Hermitian systems, we will show that in non-Hermitian systems, it can lead to the emergence of a wave-packet, similar to the effect seen earlier (caused by the reflection of the Hermitian wave-packet from the wall). However, in this case, the cause is the disorder. As we will discuss in detail below, the main reason for this behavior is that in non-Hermitian systems, the tail of the wave-packet in position space (typically insignificant in Hermitian systems) dominates the long-time dynamics due to the transformation in Eq. (6). From another perspective, weak disorder causes weak scattering to all $k$-states, but even small occupations of certain $k$-states will grow exponentially in non-Hermitian systems. This phenomenon occurs only when the initial width of the wave-packet exceeds a certain threshold, denoted as $\sigma_{c,\mathrm{dis}}$, which depends on the disorder strength and differs from the previously discussed critical value $\sigma_{c,\mathrm{ref}}$.
@@
Examining the wave-packet in $k$-space, the disorder enables transitions from the center of the wave-packet (initially positioned at $k_{0}$) to all other $k$ values. In the case of a Hermitian Hamiltonian, this scattering process continues until it reaches a saturation value determined by the disorder strength $w$ and the system size. However, since the Hamiltonian is non-Hermitian, each $k$-value also experiences exponential growth or decay. In particular, the most significant $k$-value which is $k=-\pi/2$ (since it exhibits the fastest growth, as was discussed earlier), will eventually surpass the initial Gaussian wave-packet in magnitude. This explains the transition presented in the figure.
To estimate the transition time, we develop approximations for the amplitudes of both the $k=-\pi/2$ packet and the initial Gaussian wave-packet. We start with the $k=-\pi/2$ packet. Using perturbation theory and averaging over the disorder realizations (denoted by $\left\llbracket\right\rrbracket$), we find that at small times
$\left\llbracket|c_{k}(t)|^{2}\right\rrbracket-|c_{k}(0)|^{2}\approx\frac{t^{2}}{\hbar^{2}}\frac{w^{2}}{3}\frac{1}{N}.$ (26)
At intermediate times this increase saturates, with saturation value and saturation time given approximately by
$V_{s}\approx\frac{w^{2}}{3}\frac{1}{N}\frac{1}{\mathrm{Re}\left(E_{k}-E_{k_{0}}\right)^{2}},\,t_{s}\approx\frac{\pi}{\mathrm{Re}(E_{k}-E_{k_{0}})},$ (27)
and at later times the non-Hermitian exponential growth sets in:
$|\psi(k,t)|^{2}\approx V_{s}e^{2(t_{l}-t_{r})(t-t_{s})}.$ (28)
Plugging in $k_{0}=\frac{\pi}{4}$ and $k=-\pi/2$, yields our first approximation.
For the second approximation, as we observed above, up to intermediate times the Gaussian is well described by the “continuum limit” Hamiltonian. We thus obtain
@@
where $k_{m}$ is the maximum of the wave-packet, and $b=\ln\left(\frac{t_{r}}{t_{l}}\right)\sqrt{t_{l}t_{r}}$. We have verified that both of these approximations are in a very good agreement with the numerical results (see Fig. 12 in Appendix E).
By comparing Eq. (29) and (28), we obtain a quadratic equation, and its solution readily gives the transition time, marked by the green dashed line in Fig. 6. Finally, it is worth noting that even if no disorder is considered ($w=0$), numerical artifacts may induce such transition. This is because the numerical accuracy of the computer may not be sufficient to capture the evolution of the tail of the wave-packet, effectively introducing a disorder term with amplitude related to the computer precision threshold. To overcome this (for example, to plot the disorder-less scenario that was presented in Fig. 4) one may need to increase the precision by using an arbitrary precision library (such as Advanpix [68] in Matlab).
@@ V Discussion and Conclusions
Note added. After completing this work we have become aware of the very recent preprint by He and Ozawa (Ref. [80]) on wave-packet dynamics in the Hatano-Nelson model, reporting a similar boundary-associated non-Hermitian reflection (jump). However, their analysis emphasizes the point of view of temporal Goos–Hänchen shifts at the edge, while we uncover the coexistence of the Hermitian and non-Hermitian wave-fronts as the basis for this phenomenon.
@@ Appendix D Calculation of the transition time of the non-Hermitian reflection
Keeping $a_{0}$ to be some reference lattice constant (where $a\ll a_{0}$), we have $v_{h}=2a\sqrt{t_{r}t_{l}}$ where $\sqrt{t_{l}t_{r}}=\frac{a_{0}^{2}}{a^{2}}\sqrt{t_{l,0}t_{r,0}},$ and $r=\left(\frac{t_{r,0}}{t_{l,0}}\right)^{a/2a_{0}}$ (see Appendix C). Plugging this into Eq. (81,82), we finally obtain